Suppose that a typical firm in a monopolistically competitive industry faces a demand curve given by:


q = 60 − (1/2)p, where q is quantity sold per week.


The firm’s marginal cost curve is given by: MC = 60.

  1. How much will the firm produce in the short run? 
  2. What price will it charge?

In addition to providing the quantitative answers for the question, please also describe the approach you used to arrive at your conclusions.


1
Expert's answer
2021-02-26T10:38:18-0500

Q=60 - "\frac{1}{2}P"

TR= P×Q

2Q= 120 - P

P=120 - 2Q

TR= (120 -2Q)Q

TR= 120Q - 2Q2

TR'=MR= 120 - 4Q

using the profit-maximizing approach

where MC=MR

and MC is 60 (given)


  1. 60= 120 - 4Q

60 = 4Q

Q= 15


2) P= 120 - 2Q

where Q is 15

P= 120 - 2(15)

P= 120 - 30

P= 90


The approach used is the profit maximization approach. It equates the marginal revenue (MR) to the marginal cost (MC) to arrive at the profit maximizing quantity, Q, (15) in this case and the profit maximizing price, P, (90) in for the firm.